Find each element in the circuit's time varying voltage v(t), current i(t) and power p(t).t=0 422 0000 2 mH + 15 V i(t) 20 uF ñ vc(t) ]
Answer #1

to M &m zch ROUT ཙཀ act Ang at tao, switch is closed 16-o, veloco Tucs= Usults + Vbice | Ussch), at too ما 150 (VW) = 153 Butad to a Colao 0 = 15+ A M = -15 deich floog) ebol Cancs (well)t a, ca cel Juzbach Caramel som bildet ab teet Aufonul 0 = -100.: adb= l dicb . de 7 scelt - 3.06 1 snooll) t Locob loool (isxod si e (Syd 8면 am Go.o2) loool poo) / 7348455 Shuadh -Зовухахwe already found, iach, acts, ORCH, (CE) in power Pelb= Occbice مكة . طبعا =ط© PROD=UR do it -Peuples = Lich-icho =